# An exact analytical solution for Dicke superradiance

### v3

@@ Abstract
For numerical simulations, the elements of the matrix above can be written as $\mathcal{P}_{mj}\theta_{j\geq m}\theta_{j\geq\bar{m}}+\mathcal{R}_{mj}\theta_{j\geq m}\theta_{\bar{m}\geq j}$ where only the $\mathcal{R}_{mj}$ have an explicit time dependence.
@@
Conclusions and outlook - We have provided a compact, fully analytical solution to the Dicke superradiance process, within the fully symmetric subspace of $N$ quantum emitters. The main observations have been i) the existence of simple recursive relations between Dicke states of index $m$ and $m-1$ (also previously observed in other works such as Refs. [11]), ii) the possibility to find a close analytical solution for the recursive equation and iii) the identification of this solution as a sum of residues of a given simple function. This suggests that the solution could be expressed as a complex integral on a closed contour in the complex plane, which we will investigate in the future. We have also shown that the same simple formulation can be extended to the time evolution of any initial Dicke state as an initial state of the system.

### v4

@@ Abstract
For numerical simulations, the elements of the matrix above can be written as $\mathcal{P}_{mj}\theta_{j\geq m}\theta_{j>\bar{m}}+\mathcal{R}_{mj}\theta_{j\geq m}\theta_{\bar{m}\geq j}$ where only the $\mathcal{R}_{mj}$ have an explicit time dependence.
@@
Conclusions and outlook - We have provided a compact, fully analytical solution to the Dicke superradiance process, within the fully symmetric subspace of $N$ quantum emitters. The main observations have been i) the existence of simple recursive relations between Dicke states of index $m$ and $m-1$ (also previously observed in other works such as Refs. [11]), ii) the possibility to find a closed analytical solution for the recursive equation and iii) the identification of this solution as a sum of residues of a given simple function. This suggests that the solution could be expressed as a complex integral on a closed contour in the complex plane, which we will investigate in the future. We have also shown that the same simple formulation can be extended to the time evolution of any initial Dicke state as an initial state of the system.